You sold short 100 shares of Macy's, Inc. common stock on margin at $29.78 per share. Assume the initial margin is 50% and the maintenance margin is 30%. One year later, the stock price closes at $17.00, and it has paid cash dividends of $1.51 per share. What is your return on equity? Ignore margin interest.
Answer #1

A short seller of stock gains when stock price actually decreases beyond the price at which it was sold. Here stock was sold at 29.78 per share and stock price closes at 17 per share. Hence gain = 29.78 - 17.00 = 12.78 per share.

Thus total gain = 12.78 * 100 = 1278

But, by short selling the stock, there is opportunity loss of not receiving the cash dividend of 1.51 * 100 = 151

Hence net gain = 1278 - 151 = 1127

Equity = Initial margin (or ) Investment made = 100 * 29.78 * 50% = 1489

Return on Equity = 1127 / 1489 * 100 = 75.69 %